OpenStudy (anonymous):

Question Part Points Submissions Used Your college newspaper, The Collegiate Investigator, sells for 90¢ per copy. The cost of producing x copies of an edition is given by C(x) = 50 + 0.10x + 0.001x2 dollars. (a) Calculate the marginal revenue R'(x) and profit P'(x) functions. HINT [See Example 2.]

OpenStudy (anonymous):

90¢ per copy R(x)=o.90x cost C(x) = 50 + 0.10x + 0.001x^2 dollars. C'(x)=0.1+0.002x C'(1)=0.1+0.002(1) C'(1)=$0.102 marginal cost to produce one unit Profit P(x): P(x)=R(x)-C(x) P(x)=0.90x-[50 + 0.10x + 0.001x^2] P(x) =-50+0.80x-0.001x^2 Marginal profit P'(x) P'(x)=0.80-0.002x P'(x)=0.80-0.002x0.80-0.002(1) P'(x)=$0.798 Marginal profit for the sale of one unit ..ans marginal revenue R'(x)=$0.90 ans